Twice the square of an integer is 5 less than 7 times the integer.
Answer by checkley71(8403)

2X^2=7X-5
2X^2-7X+5=0
(2X-5)(X-1)=0
2X-5=0
2X=5
X=5/2
X=2.5 ANSWER.
X-1=0
X=1 ANSWER.